INAF LBC+MODS+LUCI queue observing, 2012 October 18-19

Observer: M. Faccini, E. Sani, K. Boutsia
Telescope Operator: D. Gonzalez Huerta
Support Astronomer: D. Thompson
Instrument: MODS + LBC


Clouds in the horizon and thin scattered cirrus all over the sky. Not photometric at the beginning of the night.
We start with a MODS project that has loose constraints.
Problems with guiding and collimation with LBC.
We dectec fringing features on red channel, not explained

Data Summary

Project_RUN Objects Filter ExpTime Notes
Lardo_5B M2 SGB blue_grating 2hr15min seeing variable, ~1.0" for 2 exposures and 1.2"-1.5" for the other, probably because of passing clouds.
Giallongo_27A MS0015 g-sloan, r-sloan 1hr50min after allseeing blue:1.1" / red:1.0" - problems with guiding and collimation, fringing features in some red images
Giallongo_27B MAC0717 g-sloan, r-sloan 1hr40min after allseeing blue:1.0" / red:1.0" - problems with guiding and collimation, fringing features in some red images
Standard bd284211 blue_grating   airmass 1.1


01:27 acqMODS bd284211.acq mods1r 0002-0003
  • seeing 1.1"/1.2" on the guider
  • modsAlign -r mods1r.20121019.0002.fits

01:34 execMODS -b BlueGrating bd284211.obs mods1b 0002-0004

01:45 Starting with Lardo M2 SGB (ID551606)
  • acqMODS mods.4.m2SGBNew.acq mods1b 0005-0007
  • modsAlign mods.4.m2SGBNew.mms mods1b.20121019.0005.fits mods1b.20121019.0006.fits

02:02 execMODS mods.4.m2SGB.obs mods1b 0008-0010
  • average seeing on GCS 1.2"
  • clouds passing by during the second exposure... seeing is variable between 1.2"-1.5"

04:28 the script seems to have finished, the images are saved and displayed but the script thinks that it is still doing the 3rd exposure.
  • we abort the script and we take two confirmatory images to see if all steps work properly mods1b 0011 / mods1r 0004


4:45 authorised to use LBc

5:01 dofpia
  • problems with pointing OB it was set to TYPE=FOCUS instead of OBJECT, preset lbt rotator tracking was not ticked

5:15 Start Giallongo27A a
  • first image JUMP, stars doubled
  • second images have the same problem different direction

5:33 abort OB and run dofpia again

5:36 restart OB giallongo27A a

5:42 abort OB and stop guiding

5:43 restart OB without guiding
  • no more elongation or JUMP

6:23 dofpia
  • lbcrangebal

6:36 start OB Giallongo 27A b
  • collimation losses at the end of the each OB

7:17 dofpia
  • strange fringing

7:24 start OB Giallongo 27A c *frame 072416 has a strange pattern in the red channel that disappear in the following frame + a zoom highlight a fringing pattern
  • frame 073209 r-channel, possible transit of a satellite/plane...
  • again strange patter in the 073728 r-channel, maintained (but fainter) also in the next frame: 074009 + fringing
  • again strange patter in the 075546 r-channel, maintained (but fainter) also in the next frame

8:05 dofpia

8:09 lbcrangebal

8:13 start OB Giallongo 27A d
  • frame r082203 and b082322 plane/satellite pattern

8.32 OB aborted due to significant elongation of the stars
  • dofpia

8.48 doftpia failed to converge. Clearing the forces on the primary (red)

8.57 re-stop of dofpia and cleaning of the primary mirror optics

8.59 re-run dofpia
  • lbcrangebal

9.07 re-run MS0015.9+16_GR_d.ob

9:16 aborted cause elongation

9:17 dofpia

9:19 start again MS0015.9+16_GR_d.ob with less elongation

9.56 stop run Giallongo 27A

9:58 start pointing Giallonfgo27B MAC0717+37_GR_POINTING.OB
  • dofpia
  • lbcrangebal

10.15 start MAC0717+37_a.ob with re-inserting the GUIDING
  • frame r102903 - r103141 - r103419 - r103657 strange pattern, the first 3 also with fringing

10:56 dofpia

11:01 start MAC0717+37_b.ob

11:43 MAC0717+37_b.ob completed, few frames with little elongation * dofpia

11:47 start MAC0717+37_c.ob

12:28 end MAC0717+37_c.ob. The last frame has high background counts

12:32 still cloudy, impossible to perform flats
  • Close Dome

  • execMODS mods1b 0012-0016
  • execMODS mods1b 0017-0019 with 1.0" slit
  • execMODS mods1b 0020-0022 saturated - Abort
  • execMODS mods1b 0023-0032
  • execMODS mods1b 0033-0035 with 0.6" slit
  • execMODS mods1b 0036-0040
  • execMODS mods1b 0041-0046 ABORT - saturated! repeat!!!
  • execMODS mods1b 0047-0056 / mods1r 0005-0014

02:40 End of night

-- KonstantinaBoutsia - 19 Oct 2012
Topic revision: r10 - 28 Oct 2012, KonstantinaBoutsia
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback